OpenStudy (amypurplediaz28):

(05.01) Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system?

OpenStudy (mathmale):

You have not listed the four possible answers. Before you go any further, look up and type out the definition of "inconsistent system of linear equations."

OpenStudy (amypurplediaz28):

(05.01) Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = −6 x − 2y = 7 8x − 4y = −5

OpenStudy (mathmale):

that's better. Now, what is the definition of "inconsistent-independent system?"

OpenStudy (amypurplediaz28):

inconsistent has no solutions

OpenStudy (amypurplediaz28):

or non campatible with the other @mathmale

OpenStudy (mathmale):

And so, what is your answer to this question?

OpenStudy (amypurplediaz28):

i am not sure i think D?

OpenStudy (mathmale):

Explain your answer, please.

OpenStudy (mathmale):

You are given 4x + 2y = 9. You "think" the answer equation for which this system is inconsistent is D. D is 8x − 4y = −5. It can be shown that the system 4x + 2y = 9 8x - 4y = -5 has a solution. Therefore, D could not be correct. Choose another answer choice. Determine whether or not the given equation and your answer choice have a solution.

OpenStudy (mathmale):

You are given four possible answer choices. You must find the answer choice that produces an inconsistent system, or, in other words, produces a system that has no solution. Work through these four answer choices until you find the one that produces a system with no solution. Hint: it's NOT D.